# Efficient Vocal-Conditioned Music Generation via Soft Alignment Attention and Latent Diffusion

### v1

@@ Abstract
We present a lightweight latent diffusion model for vocal-conditioned musical accompaniment generation that addresses critical limitations in existing music AI systems. Our approach introduces a novel soft alignment attention mechanism that adaptively combines local and global temporal dependencies based on diffusion timesteps, enabling efficient capture of multi-scale musical structure. Operating in the compressed latent space of a pre-trained variational autoencoder, the model achieves a 220× parameter reduction compared to state-of-the-art systems while delivering 52× faster inference. Experimental evaluation demonstrates competitive performance with only 15M parameters, outperforming OpenAI Jukebox in production quality and content unity while maintaining reasonable musical coherence. The ultra-lightweight architecture enables real-time deployment on consumer hardware, making AI-assisted music creation accessible for interactive applications and resource-constrained environments.
@@ I Introduction
This paper introduces an ultra-lightweight latent diffusion model specifically designed for vocal-conditioned accompaniment generation. Our approach directly accepts vocal melodies as input and generates corresponding instrumental backing tracks while maintaining harmonic consistency, rhythmic alignment, and stylistic coherence. The model operates in compressed latent space using a novel soft alignment attention mechanism that dynamically balances local acoustic patterns with global musical structure.
@@ II-B Conditional Audio Generation and Attention Mechanisms
The success of transformer architectures [13] highlighted attention mechanisms’ importance for modeling long-range dependencies, though standard self-attention scales quadratically with sequence length, making it prohibitive for audio applications. This motivated efficient attention variants including local attention [14], sparse attention patterns [15], and hybrid approaches combining multiple scales.
Rotary position embeddings (RoPE) [16] provide relative position information through rotation matrices, demonstrating effectiveness in audio applications where relative temporal relationships matter more than absolute positions. Our work builds upon these foundations by introducing a soft alignment attention mechanism that adaptively combines local and global patterns based on diffusion timesteps.
@@ III Methodology
III-A Problem Formulation and Dataset
We formulate vocal-conditioned accompaniment generation as a conditional diffusion process in latent space. Given a vocal melody $v\in\mathbb{R}^{T\times F}$ with temporal frames $T$ and frequency bins $F$, we generate a corresponding accompaniment $a\in\mathbb{R}^{T\times F}$ maintaining harmonic consistency, rhythmic alignment, and stylistic coherence.
The process operates in latent space using a pre-trained VAE, mapping audio to compressed representations $z_{v},z_{a}\in\mathbb{R}^{64\times 1024}$ with 2048:1 temporal compression while preserving perceptually relevant information. This enables tractable diffusion modeling while maintaining musical structure necessary for coherent generation.
Our training dataset comprises approximately 10,000 vocal-accompaniment pairs from three sources: MUSDB18 [17] providing 150 professionally separated tracks, MoisesDB [18] contributing additional multi-track recordings, and a web-scraped collection of popular music from 2000 onward. Web-scraped data was obtained using the Song Popularity Dataset [19] for track identification, followed by automated acquisition and source separation using Demucs [20]. The complete collection process required approximately 50 hours of computational time.
All audio underwent standardized preprocessing: conversion to 44.1 kHz sampling rate, peak normalization, and segmentation into 47-second chunks. The 47-second duration aligns with Stable Audio Open architecture, containing exactly 2,097,152 samples at 44.1 kHz, which with 2048:1 compression resolves to 1024 tokens—computationally convenient for efficient attention computation and memory utilization.
@@ III-B Latent Space Encoding
We employ the VAE component from Stable Audio Open [4], representing state-of-the-art audio compression optimized for generative modeling. The encoder architecture builds on the Descript Audio Codec framework, featuring fully-convolutional design enabling arbitrary-length processing while maintaining high reconstruction fidelity.
The encoder bottleneck is parameterized as a variational autoencoder with 64-dimensional latent space, providing continuous representation more suitable for diffusion modeling than discrete tokenization. The decoder employs transposed strided convolutions with symmetric channel contraction and weight normalization throughout. Critically, the decoder omits traditional $\tanh$ activation at output, as empirical evaluation revealed this nonlinearity introduced harmonic distortion artifacts detrimental to musical content quality.
This VAE architecture achieves 2048:1 compression ratio, mapping 47-second clips to 1024-dimensional latent sequences while preserving musical structure and timbral characteristics. The combination of dilated convolutions, Snake activations [21], and careful architectural choices captures both local acoustic patterns and global musical structure, providing optimal foundation for vocal-conditioned diffusion.
@@ IV Model Architecture
IV-A U-Net Design with Soft Alignment Attention
Our model employs a U-Net architecture enhanced with novel attention mechanisms specifically designed for vocal-accompaniment relationships. The network processes latent representations $(B,64,1024)$ through hierarchical encoder-decoder structure featuring multiple resolution levels, skip connections, and sophisticated conditioning mechanisms emphasizing dual-stream vocal conditioning.
The U-Net comprises downsampling blocks, bottleneck attention layer, and upsampling blocks with residual connections. Each downsampling block includes Group Normalization, SiLU activation, Feature-wise Linear Modulation (FiLM) [22], and Conv1d layers progressively reducing temporal resolution while increasing channel depth: (64,1024)(128,512)(256,256)(512,128)(64,1024$\rightarrow$(128,512$\rightarrow$(256,256$\rightarrow$(512,128)( 64 , 1024 ) → ( 128 , 512 ) → ( 256 , 256 ) → ( 512 , 128 ).
@@ IV-A U-Net Design with Soft Alignment Attention
IV-B Soft Alignment Attention Mechanism
Our soft alignment attention combines local and global attention patterns through time-dependent weighting reflecting the hierarchical diffusion process. Local attention operates within sliding windows of size 16, capturing fine-grained temporal dependencies through convolutional processing. Global attention computes relationships across entire sequences using rotary position embeddings [16] for improved positional encoding.
@@ IV-B Soft Alignment Attention Mechanism
Figure 1: Detailed soft alignment attention mechanism computing both local and global attention patterns. The dual-scale approach captures fine-grained temporal dependencies within local windows while maintaining global musical coherence through full-sequence attention, critical for maintaining musical structure and quality.
The multi-scale attention design enables the model to capture information at multiple temporal scales essential for musical coherence and quality. Local attention windows preserve fine-grained acoustic relationships between neighboring temporal frames, while global attention maintains awareness of long-range harmonic progressions and structural elements. This dual-scale approach proves particularly effective for musical applications where both micro-timing relationships and macro-structural coherence are critical for perceptual quality.
@@ IV-C Architecture Specifications and Training Objective
The complete model comprises approximately 15 million parameters distributed across U-Net backbone, attention mechanisms, and conditioning networks. Attention mechanisms utilize 8 heads with 64-dimensional projections, and local attention window size 16 provides effective local modeling while maintaining computational tractability.
The model employs v-parametrization [23] rather than standard noise prediction, providing improved training stability and sample quality for audio applications:
@@ V Training and Implementation
The model training employed AdamW optimization with learning rate $3.5\times 10^{-4}$ and cosine annealing schedule. Batch size selection proved critical for training stability—systematic experimentation revealed batch size 16 provided optimal balance between stable gradient estimation and computational efficiency. This stems from interactions between group normalization layers and signal noise ratio weighted loss function during diffusion.
@@
The diffusion process utilized cosine noise schedule with 800 timesteps, and classifier-free guidance training incorporated 10% conditioning dropout probability. A 10% held-out validation set monitored training progress with early stopping based on validation loss convergence. Mixed precision training provided computational efficiency while maintaining numerical stability.
@@ VI-A Evaluation Methodology and Quantitative Results
We evaluated generated accompaniments using AudioBox-Aesthetics [12], providing objective metrics across four dimensions: Production Quality (PQ) measuring technical fidelity, Production Complexity (PC) assessing arrangement sophistication, Content Enjoyment (CE) evaluating subjective appeal, and Content Unity (CU) measuring musical coherence. Evaluation compared our model against Stable Audio Open, MusicGen, OpenAI Jukebox, and ground truth across 100 generated samples and 50 reference tracks.
@@
OpenAI Jukebox 5.0B $>$6h Lyrics + Genre
@@
Our Model 15M 21s Vocal Melody
Results reveal competitive performance considering our model’s smaller scale and training constraints. While achieving lower absolute scores compared to larger systems, our model outperforms OpenAI Jukebox in Production Quality (5.73 vs 5.30) and Content Unity (4.97 vs 4.66), attributed to Jukebox’s tendency toward noise artifacts. The Production Complexity score (5.14) demonstrates that despite resource constraints, soft alignment attention enables reasonably sophisticated musical arrangements through direct vocal conditioning.
Our model achieves exceptional efficiency with only 15M parameters—220× smaller than MusicGen Large, 20× smaller than MusicGen Small, 80× smaller than Stable Audio Open, and 333× smaller than OpenAI Jukebox. Inference speed of 21 seconds per 47-second audio represents substantial advantages: 52× faster than MusicGen Large and 11× faster than MusicGen Small. While Stable Audio Open achieves comparable inference time, it cannot process direct musical input, positioning our model as optimal for audio-to-audio conditioning.
@@ VI-B Qualitative Analysis and Performance Trade-offs
A notable limitation is the model’s current inability to generate distinct instrumental sounds with clear acoustic characteristics, reflected in lower content enjoyment scores compared to models trained on larger datasets. Generated accompaniments lack timbral diversity typical of professional recordings, producing more homogeneous sonic palettes dominated by synthesized textures. This limitation stems from compressed latent representation, limited training data, and focus on overall musical structure rather than fine-grained timbral detail—constraints inherent to ultra-lightweight design philosophy.
@@ VII Discussion and Future Work
The presented vocal-conditioned diffusion model demonstrates direct musical input viability for accompaniment generation, addressing a significant gap in current music AI systems. The soft alignment attention mechanism provides effective solutions for capturing multi-scale temporal dependencies in audio, while v-parametrization ensures stable training dynamics suitable for complex musical content.
@@ VII-A Efficiency Advantages and Deployment Considerations
Performance trade-offs inherent in efficiency gains are evident in AudioBox-Aesthetics evaluation, where our model achieves lower absolute scores compared to larger systems. However, the ability to outperform OpenAI Jukebox demonstrates that architectural innovation can compensate for scale limitations. Relatively modest scores in content enjoyment and unity reflect challenges of training with fewer parameters and limited data, yet the model successfully synthesizes complete multi-track accompaniments—remarkable given resource constraints.
@@ VII-B Training Insights and Limitations
Experimentation revealed important phenomena regarding model architecture and training dynamics. A critical finding concerns the relationship between model complexity and perceptual quality: increasing U-Net depth consistently reduced training loss but did not correlate with improved audio quality. This suggests traditional loss metrics may inadequately capture perceptual aspects of musical generation, highlighting fundamental challenges where mathematical optimization objectives diverge from human preferences.
@@
Future work should explore hierarchical generation approaches handling multiple temporal scales simultaneously, enabling coherent composition of complete songs. Variable-length audio processing could increase training data availability and reduce computational requirements. Multi-modal conditioning combining vocal, textual, and symbolic inputs could provide enhanced control for professional workflows. Real-time optimization would enable live performance accompaniment, while style-specific training could improve generation quality for particular genres.

### v2

# SAMUeL: Efficient Vocal-Conditioned Music Generation via Soft Alignment Attention and Latent Diffusion

@@ Abstract
We present SAMUeL (Soft Alignment Music U-net Latent diffusion) [^1] Publicly available at: https://github.com/HaysonC/SAMUeL-GEN, a lightweight latent diffusion model for vocal-conditioned musical accompaniment generation that addresses critical limitations in existing music AI systems. SAMUeL introduces a novel soft alignment attention mechanism that adaptively combines local and global temporal dependencies based on diffusion timesteps, enabling efficient capture of multi-scale musical structure. Operating in the compressed latent space of a pre-trained variational autoencoder, SAMUeL achieves a 220× parameter reduction compared to state-of-the-art systems while delivering 52× faster inference. Experimental evaluation demonstrates competitive performance with only 15M parameters, outperforming OpenAI Jukebox in production quality and content unity while maintaining reasonable musical coherence. The ultra-lightweight architecture enables real-time deployment on consumer hardware, making AI-assisted music creation accessible for interactive applications and resource-constrained environments.
@@ I Introduction
This paper introduces SAMUeL (Soft Alignment Music U-net Latent diffusion), an ultra-lightweight latent diffusion model specifically designed for vocal-conditioned accompaniment generation. The model name reflects its core architectural components: the Soft Alignment attention mechanism that dynamically balances temporal dependencies, the U-net backbone architecture, and operation in compressed latent space for efficiency. SAMUeL directly accepts vocal melodies as input and generates corresponding instrumental backing tracks while maintaining harmonic consistency, rhythmic alignment, and stylistic coherence. SAMUeL operates in compressed latent space using a novel soft alignment attention mechanism that dynamically balances local acoustic patterns with global musical structure.
@@ II-B Conditional Audio Generation and Attention Mechanisms
Transformer vs. Ours SAMUeL employs a U-Net architecture augmented with attention mechanisms inspired by transformers [13] but is distinct in its design. Unlike standard transformers, which rely heavily on Multi-Layer Perceptrons (MLPs) and Layer Normalization contributing significantly to parameter count, SAMUeL omits these components to achieve an ultra-lightweight architecture. Our soft alignment attention mechanism, combining local and global patterns, is tailored for audio applications, leveraging rotary position embeddings (RoPE) [14] for global attention while maintaining computational efficiency.
@@ III Methodology
III-A Problem Formulation
We formulate vocal-conditioned accompaniment generation as a conditional diffusion process in latent space. Given a vocal melody $v\in\mathbb{R}^{2\times T}$ with temporal frames $T$ and 2 channels (stereo audio), we generate a corresponding accompaniment $a\in\mathbb{R}^{2\times T}$ maintaining harmonic consistency, rhythmic alignment, and stylistic coherence.
The process operates in latent space using a pre-trained music variational autoencoder (VAE), mapping audio to compressed representations $z_{v},z_{a}\in\mathbb{R}^{64\times T/r}$ with $r$ being the compression rate of the VAE. This enables tractable diffusion modeling while maintaining musical structure necessary for coherent generation.
@@ III-B Latent Space Encoding
We employ the pre-trained VAE component from Stable Audio Open [4], representing state-of-the-art audio compression optimized for latent diffusion generation. The encoder architecture builds on the Descript Audio Codec framework[15], featuring fully-convolutional design enabling arbitrary-length processing while maintaining high reconstruction fidelity.
The encoder bottleneck is parameterized as a variational autoencoder with 64-dimensional latent space and a compression rate, $r=2048$, providing continuous representation more suitable for diffusion modeling than discrete tokenization. The decoder employs transposed strided convolutions with symmetric channel contraction and weight normalization throughout. Critically, the decoder omits traditional $\tanh$ activation at output, as empirical evaluation revealed this nonlinearity introduced harmonic distortion artifacts detrimental to musical content quality.
III-C Data Collection
For this study, we introduce POPDB [16], a dataset of 15,713 vocal-accompaniment pairs derived from popular songs spanning 2000 onward, totaling approximately 200 hours of audio. The dataset construction leveraged the Song Popularity Dataset [17] from Kaggle, which catalogs popular songs with metadata, selected for its diversity across genres and eras.
Audio was acquired using the yt-dlp Python library for automated YouTube downloading, searching by song titles and filtering out playlists and content exceeding 15 minutes. Audio was processed in WAV format at 44.1 kHz to maintain fidelity. Source separation used Demucs [18] to isolate vocal and accompaniment components. Each song was encoded using the Stable Audio Open VAE, producing 1024-token sequences corresponding to 47.55 seconds of audio, then chunked sequentially without overlap. Trailing chunks shorter than 47.55 seconds were discarded.
The 47.55-second chunk size aligns with Stable Audio Open’s VAE optimization, balancing efficiency and performance. We hypothesize that at least half of each chunk contains vocals, based on typical song structures where non-vocal segments (e.g., intros, bridges) are usually under 20 seconds [19]. This assumption requires validation in future work to ensure robust vocal conditioning. The pipeline required approximately 50 hours of computational time, resulting in POPDB containing 15,713 encoded vocal-accompaniment pairs for training.
@@ IV Model Architecture
IV-A U-Net Design
SAMUeL employs a U-Net architecture enhanced with novel attention mechanisms specifically designed for vocal-accompaniment relationships. The network processes latent representations $\left(B,64,1024\right)$ through a hierarchical encoder-decoder structure featuring multiple resolution levels, residual connections, skip connections, and sophisticated conditioning mechanisms emphasizing dual-stream vocal conditioning.
The U-Net comprises downsampling blocks, a bottleneck attention layer, and upsampling blocks with residual connections. Each downsampling block includes Group Normalization, SiLU activation, Feature-wise Linear Modulation (FiLM) [20], and Conv1D layers, progressively reducing temporal resolution while increasing channel depth:
$(64,1024)\rightarrow(128,512)\rightarrow(256,256)\rightarrow(512,128).$
SiLU (Sigmoid Linear Unit) [21], also known as Swish, is chosen for its smooth, non-monotonic activation properties, which improve gradient flow and yield better empirical performance in deep sequence models.
This architectural design allows the network to efficiently capture both fine-grained local dependencies and long-range structure in temporal data. By leveraging FiLM and attention-based soft alignment, SAMUeL dynamically integrates conditioning information at multiple scales, enabling more expressive modeling of vocal-guided accompaniment generation.
IV-B Soft Alignment Attention Mechanism
@@
Our soft alignment attention combines local and global attention patterns through time-dependent weighting reflecting the hierarchical diffusion process. Local attention operates within sliding windows of size 16 on the temporal axis; every chunk is padded with zeros at the beginning and the end. While local attention does not leverage positional embedding to focus on the musical harmonic structure where absolute position is less important, global attention computes relationships across entire sequences using rotary position embeddings [14] for improved positional encoding.
The underlying intuition is to expose the model to the full structure of the song early on, while progressively shifting focus toward fine-grained acoustic interactions between nearby temporal frames as diffusion proceeds. This dual-scale attention strategy allows SAMUeL to model both the precise timing details and the broader musical form, which are critical for generating coherent and expressive musical accompaniments.
@@
Figure 1: Detailed soft alignment attention mechanism computing both local and global attention patterns. The dual-scale approach captures fine-grained temporal dependencies within local windows while maintaining global musical coherence through full-sequence attention, critical for maintaining musical structure and quality. Note that RoPE is not applied to for local attention, learning about the general local acoustic relationships.
The multi-scale attention design enables SAMUeL to capture information at multiple temporal scales essential for musical coherence and quality. Local attention windows preserve fine-grained acoustic relationships between neighboring temporal frames, while global attention maintains awareness of long-range harmonic progressions and structural elements. This dual-scale approach proves particularly effective for musical applications where both micro-timing relationships and macro-structural coherence are critical for perceptual quality.
@@ IV-C Architecture Specifications and Training Objective
SAMUeL comprises approximately 15 million parameters distributed across U-Net backbone, attention mechanisms, and conditioning networks. The U-Net backbone contains 3 downsampling and 3 upsampling convolutional blocks, each having two convolutional layers and one FiLM layer. Attention mechanisms utilize 8 heads with 64-dimensional projections, and local attention window size 16 provides effective local modeling while maintaining computational tractability.
SAMUeL employs v-parametrization [22] rather than standard noise prediction, providing improved training stability and sample quality for audio applications:
@@ V Training and Implementation
SAMUeL training employed AdamW optimization [23] with learning rate $3.5\times 10^{-4}$ and cosine annealing schedule for one hour on one A100. Batch size selection proved critical for training stability—systematic experimentation revealed batch size 16 provided optimal balance between stable gradient estimation and computational efficiency. This stems from interactions between group normalization layers and signal noise ratio weighted loss function during diffusion.
@@
[…]
@@ VI-A Evaluation Methodology and Quantitative Results
We evaluated generated accompaniments using AudioBox-Aesthetics [12], a pre-trained transformer-based model developed by Meta that provides comprehensive audio quality assessment. This model addresses critical limitations of existing audio evaluation metrics such as Fréchet Audio Distance (FAD), which produces only overall scores without capturing the nuanced aspects of musical quality. Unlike traditional benchmarks that provide uninformative evaluations and amplify individual bias from human-annotated data, AudioBox-Aesthetics represents the only available open-source pre-trained model capable of multi-dimensional audio evaluation.
The AudioBox-Aesthetics model provides objective metrics across four independent dimensions, each scored on a scale from 1 (less desirable) to 10 (more desirable): Production Quality (PQ) measuring technical fidelity and recording quality, Production Complexity (PC) assessing arrangement sophistication and musical intricacy, Content Enjoyment (CE) evaluating subjective appeal and listener satisfaction, and Content Unity (CU) measuring musical coherence and structural consistency. The model was trained on 500 hours of audio samples annotated by 158 expert raters, ensuring robust evaluation standards. The training process emphasized category independence, with low correlation between evaluation dimensions demonstrating that each metric captures distinct aspects of audio quality. Rating distributions span the full evaluation scale, indicating diversity in the collected training data and comprehensive coverage of quality variations.
The original validation study confirmed high correlation between the model’s predicted quality scores and ground truth labels from expert human evaluators, establishing its reliability for objective music quality assessment. This multi-dimensional approach proves particularly valuable for evaluating generative music models, where different aspects of quality may vary independently based on architectural choices and training constraints.
Evaluation compared SAMUeL against Stable Audio Open, MusicGen, OpenAI Jukebox, and ground truth across 10 generated samples, and the average score of the ten tracks are reported in Table I. Due to the small sample size, results are preliminary and require larger-scale validation in future work.
@@
OpenAI Jukebox 5.0B >>6h1 Lyrics + Genre
@@
SAMUeL 15M 21s Vocal Melody
  • 1
    Result taken from [7]
Results reveal competitive performance considering SAMUeL’s smaller scale and training constraints. While achieving lower absolute scores compared to larger systems, SAMUeL outperforms OpenAI Jukebox in Production Quality (5.73 vs 5.30) and Content Unity (4.97 vs 4.66), attributed to Jukebox’s tendency toward noise artifacts. The Production Complexity score (5.14) demonstrates that despite resource constraints, soft alignment attention enables reasonably sophisticated musical arrangements through direct vocal conditioning.
SAMUeL achieves exceptional efficiency with only 15M parameters—220× smaller than MusicGen Large, 20× smaller than MusicGen Small, 80× smaller than Stable Audio Open, and 333× smaller than OpenAI Jukebox. Inference speed of 21 seconds per 47-second audio represents substantial advantages: 52× faster than MusicGen Large and 11× faster than MusicGen Small. While Stable Audio Open achieves comparable inference time, it cannot process direct musical input, positioning SAMUeL as optimal for audio-to-audio conditioning.
@@ VI-B Qualitative Analysis and Performance Trade-offs
A notable limitation is SAMUeL’s current inability to generate distinct instrumental sounds with clear acoustic characteristics, reflected in lower content enjoyment scores compared to models trained on larger datasets. Generated accompaniments lack timbral diversity typical of professional recordings, producing more homogeneous sonic palettes dominated by synthesized textures. This limitation stems from compressed latent representation, limited training data, and focus on overall musical structure rather than fine-grained timbral detail—constraints inherent to ultra-lightweight design philosophy.
@@ VII Discussion and Future Work
SAMUeL demonstrates direct musical input viability for accompaniment generation, addressing a significant gap in current music AI systems. The soft alignment attention mechanism provides effective solutions for capturing multi-scale temporal dependencies in audio, while v-parametrization ensures stable training dynamics suitable for complex musical content.
@@ VII-A Efficiency Advantages and Deployment Considerations
Performance trade-offs inherent in efficiency gains are evident in AudioBox-Aesthetics evaluation, where SAMUeL achieves lower absolute scores compared to larger systems. However, the ability to outperform OpenAI Jukebox demonstrates that architectural innovation can compensate for scale limitations. Relatively modest scores in content enjoyment and unity reflect challenges of training with fewer parameters and limited data, yet SAMUeL successfully synthesizes complete multi-track accompaniments—remarkable given resource constraints.
@@ VII-B Training Insights and Limitations
Experimentation revealed important phenomena regarding model architecture and training dynamics. A critical finding concerns the relationship between model complexity and perceptual quality: increasing U-Net depth consistently marginally reduced training loss but did not correlate with improved audio quality. This suggests traditional loss metrics may inadequately capture perceptual aspects of musical generation, highlighting fundamental challenges where mathematical optimization objectives diverge from human preferences. Further works could explore alternative training objectives that better align with perceptual quality, such as adversarial training or perceptual loss functions.
@@
A significant constraint affecting SAMUeL’s performance stems from limited training data due to computational resource restrictions. Our model was trained on approximately 200 hours of audio data, substantially less than the 7,300 hours utilized by Stable Audio Open [4]. This 36-fold disparity in training data represents a critical limiting factor, as deep generative models typically exhibit strong scaling behavior with increased data availability. Despite this constraint, SAMUeL demonstrates competitive performance relative to its training data scale, suggesting that the sophisticated soft alignment attention architecture effectively leverages available information. We anticipate that scaling to comparable dataset sizes would yield substantial performance improvements, particularly in timbral diversity and content enjoyment metrics where data-driven learning is most beneficial.
Future work should explore hierarchical generation approaches handling multiple temporal scales simultaneously, enabling coherent composition of complete songs. Variable-length audio processing could increase training data availability and reduce computational requirements. Multi-modal conditioning combining vocal, textual, and symbolic inputs could provide enhanced control for professional workflows.